# Two-dimensional Entanglement-assisted Quantum Quasi-cyclic Low-density Parity-check Codes

### v1

@@ Abstract
For any positive integer $g\geq 2$, we derive general conditions for the existence of a $2g$-cycle in the Tanner graph of two-dimensional ($2$-D) classical quasi-cyclic (QC) low-density parity-check (LDPC) codes. Based on these conditions, we construct a family of $2$-D classical QC-LDPC codes with girth greater than $4$ by stacking $p\times p\times p$ tensors, where $p$ is an odd prime. Furthermore, for composite values of $p$, we propose two additional families of $2$-D classical LDPC codes obtained via similar tensor stacking. In this case, one family achieves girth greater than $4$, while the other attains girth greater than $6$. All the proposed $2$-D classical QC-LDPC codes exhibit an erasure correction capability of at least $p\times p$. Based on the constructed classical $2$-D QC-LDPC codes, we derive two families of $2$-D entanglement-assisted (EA) quantum low-density parity-check (QLDPC) codes. The first family of $2$-D EA-QLDPC codes is obtained from a pair of binary $2$-D classical LDPC codes and is designed such that the unassisted part of the Tanner graph of the resulting EA-QLDPC code is free of cycles of length four, while requiring only a single ebit to be shared across the quantum transceiver. The second family is constructed from a single $2$-D classical LDPC code whose Tanner graph is free from $4$-cycles. Moreover, the constructed EA-QLDPC codes inherit an erasure correction capability of $p\times p$, as the underlying classical codes possess the same erasure correction property.
@@ I Introduction
Following the success of one-dimensional (1-D) LDPC codes in classical communication and storage systems, higher-dimensional LDPC codes have also attracted significant research interest for the design of more powerful coding schemes. This growing interest is primarily driven by the stringent reliability requirements of advanced storage technologies such as shingled magnetic recording, two-dimensional magnetic recording (TDMR) [16], and three-dimensional NAND flash memories [9], etc. In particular, a fundamental challenge in TDMR systems is that the dominant error mechanisms—including random errors, burst errors, and erasures—are inherently two-dimensional, arising from inter-track interference and inter-symbol interference. As a result, conventional $1$-D coding and decoding approaches are often inadequate and inefficient, motivating the development of native two-dimensional signal-processing and coding techniques attuned to the underlying channel structure [6, 3].
@@
2-D burst errors and erasures can be handled efficiently using $2$-D LDPC codes, which naturally extend the sparse graphical structure of classical LDPC codes to higher dimensions. In this direction, several works have investigated the construction, encoding, and decoding of $2$-D LDPC codes, demonstrating their effectiveness in mitigating structured two-dimensional error patterns [12, 8, 2]. These developments motivate the design of novel 2-D classical LDPC codes as a foundation for constructing quantum LDPC codes capable of correcting clustered errors and erasures under Markovian constraints, which are particularly relevant for quantum storage and quantum communication systems [4, 7]. Although some studies have explored 2-D quantum LDPC code constructions, important limitations remain. For example, the authors of [17] construct quantum LDPC codes from 2-D spatially coupled LDPC codes which are not tailored to mitigate 2-D burst errors and erasures. Other recent works, such as [1] and [13], focus on quantum LDPC code design without explicitly addressing the correction of 2-D burst error/erasure patterns. Consequently, existing approaches do not fully account for correlated error structures arising from Markovian effects. This gap motivates our work, in which we first construct novel 2-D classical QC-LDPC codes and then utilize them to develop 2-D quantum LDPC codes specifically designed to correct clustered errors and erasures under realistic Markovian noise models.
@@
In this work, we focus on the construction of 2-D entanglement-assisted (EA) quantum LDPC codes. We adopt the entanglement-assisted framework rather than the entanglement-unassisted setting to circumvent unavoidable cycles of length four that severely degrades the decoding performance. In other words, in the EA setting, the unassisted portion of the overall Tanner graph can be designed to be free of 4-cycles. Moreover, the entanglement-unassisted framework requires the underlying classical codes to satisfy the restrictive dual-containment constraint. In contrast, the availability of pre-shared ebits in the EA framework relaxes the stringent orthogonality requirements on the parity-check matrices, thereby allowing the construction of 2-D quantum LDPC codes from arbitrary classical codes. With the drive towards implementation of high-fidelity controlled quantum gates and caching of EPR pairs generated through a physical process, the EA construction of such 2D QLDPC codes would be a promising direction. The main contributions of this work are summarized as follows.
@@
  1. (1) We first derive a general condition for the existence of $2g$-cycles in the Tanner graphs of 2-D classical LDPC codes. We then construct several families of 2-D classical QC-LDPC codes. One family, with an exact code rate and girth greater than $4$, is obtained by stacking $p\times p\times p$ tensors, where $p$ is an odd prime. In addition, two further families are constructed by stacking $p\times p\times p$ tensors with composite $p$: one family achieves girth greater than $4$, while the other attains girth greater than $6$. Finally, we demonstrate that all the proposed code families achieve a burst erasure correction capability of at least $p\times p$.
@@
  2. (2) We then employ the constructed $2$-D classical LDPC codes to develop two families of entanglement-assisted quantum LDPC codes. In the first family, two distinct $2$-D classical LDPC codes are used to ensure that the unassisted portion of the overall Tanner graph is free of $4$-cycles, while requiring minimally a single ebit. The second family is constructed using a single $2$-D classical LDPC code, where the underlying classical code itself is free of $4$-cycles.
@@
This article is organized as follows. In Section II, we first derive a general condition for the existence of $2g$-cycles in the Tanner graphs of 2-D classical LDPC codes. We then construct several families of 2-D classical QC-LDPC codes with burst erasure correction capabilities of at least $p\times p$. In Section III, we construct two families of 2-D entanglement-assisted quantum LDPC codes: one obtained from a pair of distinct $2$-D classical LDPC codes and the other constructed using a single $2$-D classical LDPC code. Finally, Section IV concludes the paper.
II 2-D Classical LDPC Code Construction
@@
From (3), it is evident that the application of $P$, $Q$ and $R$ to the identity tensor $I_{3\text{-}D}$ results in cyclic shifts along the $x$-, $y$-, and $z$-axes, respectively. Moreover, we refer to the shifts along the $x$ -, $y$-, and $z$-axes as the $X$-, $Y$-, and $Z$-shifts, respectively. These shifts have been illustrated for the identity tensor $I_{3\text{-}D}$ of size $3\times 3\times 3$ in Fig. 1.
@@
In light of Lemma 2, we construct 2-D classical LDPC codes by choosing permutations $P$ and $Q$ that implement circular shifts along the $x$- and $y$-axes, respectively. Based on this construction, we introduce the following assumptions and terminology to define the parity-check tensor $H_{2\text{-}D}$, which is obtained as a stack of tensors corresponding to the 2-D classical LDPC codes.
@@
    We introduce a new set of coordinate axes, denoted by $i$, $j$ and $k$, aligned with the $x$-, $y$-, and $z$-directions, respectively. The indices $(x,y,z)$ are used to label the positions of individual small cubes within the tensor, while the indices $(i,j,k)$ refer to the corresponding coordinates of the tensor as a whole. This distinction allows us to clearly differentiate between local indexing within the tensor and global indexing of the tensor structure.
@@
    For given size of identity tensor, i.e., $p\times p\times p$, we first define the shifts of a tensor located at an arbitrary position $(i,j,k)$ along the $i$- and $j$-directions, which correspond to shifts along the $x$- and $y$-directions. These shifts are denoted by $a(i,j,k)$ and $b(i,j,k)$, respectively, and are functions of the indices $i$, $j$, and $k$.
@@
    For given positive integers $c$, $b$ and $h$, tensors with the prescribed shifts are stacked along the $i$-, $j$-, and $k$-directions. Within this stack, the $(i,j,k)$-th tensor, with indices satisfying $0\leq i\leq c$, $0\leq j\leq b$, and $0\leq k\leq h$, is constructed as
@@
To analyze the existence of cycles in the Tanner graph of $H_{2\text{-}D}$, we unfold the 3-D parity-check tensor $H_{2\text{-}D}$ into a 2-D parity-check matrix $H_{1\text{-}D}$ of size $ph\times bcp^{2}$. The unfolding is performed layer by layer along the vertical dimension: the first $cp$ columns of $H_{1\text{-}D}$ correspond to the first vertical layer of $H_{2\text{-}D}$, the next $cp$ columns correspond to the second vertical layer, and so on. Mathematically, the parity-check tensor $H_{2\text{-}D}$ is unfolded into the matrix $H_{1\text{-}D}$ such that the $(i,j)$-th entry of $H_{1\text{-}D}$ is the $(j\pmod{p},\left\lfloor\frac{j}{p}\right\rfloor,i)$-th entry in $H_{2\text{-}D}$.
Now, consider the closed path of length $2g$ in the parity-check tensor $H_{2\text{-}D}$ described under assumption (9). This leads to the following theorem, which establishes the existence of cycles of length $2g$ in the Tanner graph associated with $H_{2\text{-}D}$ along the corresponding closed path.
@@
Let $H_{2\text{-}D}$ be the parity-check tensor composed of permutation tensors of size $p\times p\times p$. Assume that the permutation tensor located at position $(i,j,k)$ is specified by the shift functions $a(i,j,k)$ and $b(i,j,k)$. Then the Tanner graph of $H_{2\text{-}D}$ contains a cycle of length $2g$ if and only if
$\sum_{m=0}^{g-1}\Bigl(a(i_{m},j_{m},k_{m})-a(i_{m+1},j_{m+1},k_{m})\Bigr)$ $=0\pmod{p}$,
$\sum_{m=0}^{g-1}\Bigl(b(i_{m},j_{m},k_{m})-b(i_{m+1},j_{m+1},k_{m})\Bigr)$ $=0\pmod{p}$,
@@
Throughout the proof, $(i,j,k)$ denote the indices of the tensor, whereas $(x,y,z)$ denote the local indices of the small cubes, i.e., cubes within the tensor. Consider the following closed path
@@
where $(i_{\ell},j_{\ell})\neq(i_{\ell+1},j_{\ell+1})$ and $k_{\ell}\neq k_{\ell+1}$ for $\ell=0,1,2,\ldots,g-2$, and $(i_{g-1},j_{g-1})\neq(i_{0},j_{0})$ with $k_{g-1}\neq k_{0}$. Next, consider the following sequence of indices corresponding to the nonzero entries along the given closed path:
@@
where $0\leq x_{p},y_{q},z_{r}\leq(p-1)$ under modulo $p$ for all $p,q,r$. From this point onward, we write $0\leq x,y,z\leq p-1$ without explicitly indicating $\pmod{p}$. It is straightforward to observe that, for any tensor located at position $(i,j,k)$, the $x$ ($0\leq x\leq(p-1)$) coordinate of a nonzero entry in the $z$-th layer of the tensor, where $0\leq z\leq p-1$, is given by
@@
Similarly, for any tensor located at position $(i,j,k)$, the $y$-coordinate ($0\leq y\leq p-1$) of a nonzero entry in the $z$-th layer of the tensor, where $0\leq z\leq p-1$, is given by
@@
Thus, for the existence of a $2g$-cycle corresponding to the given closed path, it suffices to show that $(x_{g},y_{g})=(x_{0},y_{0})$. Here, $(x_{g},y_{g})$ denotes the $(x,y)$-coordinates of a nonzero entry in the layer $z=z_{g-1}$ of the tensor at position $(i_{0},j_{0},k_{g-1})$, while $(x_{0},y_{0})$ denotes the $(x,y)$-coordinates of a nonzero entry in the layer $z=z_{0}$ of the tensor corresponding to $(i_{0},j_{0},k_{0})$. Now, by (4), for each index along the closed path, the $x$-coordinates of the nonzero entries are given by the following set of equations:
$x_{0}$ $=z_{0}+a(i_{0},j_{0},k_{0})\pmod{p}$,
$x_{1}$ $=z_{0}+a(i_{1},j_{1},k_{0})\pmod{p}$,
$x_{1}$ $=z_{1}+a(i_{1},j_{1},k_{1})\pmod{p}$,
$x_{2}$ $=z_{1}+a(i_{2},j_{2},k_{1})\pmod{p}$,
@@
$x_{g-1}$ $=z_{g-1}+a(i_{g-1},j_{g-1},k_{g-1})\pmod{p}$,
$x_{g}$ $=z_{g-1}+a(i_{0},j_{0},k_{g-1})\pmod{p}$.
By the above set of equations, we have
$x_{g}=x_{0}+\sum_{m=0}^{g-1}\Bigl(a(i_{m},j_{m},k_{m})-a(i_{m+1},j_{m+1},k_{m})\Bigr),$ (6)
where $(i_{g},j_{g})=(i_{0},j_{0})$. Now, by applying the same reasoning to the $y$-coordinates of the nonzero entries, we obtain that
$y_{g}=y_{0}+\sum_{m=0}^{g-1}\Bigl(b(i_{m},j_{m},k_{m})-b(i_{m+1},j_{m+1},k_{m})\Bigr),$ (7)
where $(i_{g},j_{g})=(i_{0},j_{0})$. For the existence of a $2g$-cycle, it is necessary that $(x_{g},y_{g})=(x_{0},y_{0})$. By (6) and (7), this occurs if and only if the following two equations are simultaneously satisfied:
$\sum_{m=0}^{g-1}\Bigl(a(i_{m},j_{m},k_{m})-a(i_{m+1},j_{m+1},k_{m})\Bigr)$ $=0\pmod{p}$,
$\sum_{m=0}^{g-1}\Bigl(b(i_{m},j_{m},k_{m})-b(i_{m+1},j_{m+1},k_{m})\Bigr)$ $=0\pmod{p}$,
@@
Henceforth, this section is organized into two subsections. The first subsection focuses on permutation tensors of size $p\times p\times p$ with $p$ an odd prime, while the second subsection addresses the case where $p$ is composite.
II-A 2-D Classical QC-LDPC Codes by Stacking $p\times p\times p$ Permutation Tensors ($p$ Odd Prime)
In this subsection, we focus on the construction of 2-D classical QC-LDPC codes by stacking permutation tensors of size $p\times p\times p$, where $p$ is an odd prime, while ensuring that the corresponding Tanner graph has girth greater than $4$.
Let $p$ be an odd prime, and it will carry the same meaning throughout this subsection. We construct the $2$-D QC-LDPC parity-check tensor $H_{2\text{-}D}$ by stacking $p\times p\times p^{2}$ permutation tensors, each of size $p\times p\times p$, along the $i$-, $j$-, and $k$-directions, respectively. These tensors are generated by applying various combinations of the shift operators $P$ and $Q$ to the identity tensor $I_{3\text{-}D}$, corresponding to $I$-shifts ($X$-shift) and $J$-shifts ($Y$-shifts), respectively. While constructing $H_{2\text{-}D}$, the $(i,j,k)$th permutation tensor is given by $P^{a(i,j,k)}\circ Q^{b(i,j,k)}(I_{3\text{-}D})$, where $a(i,j,k)$ and $b(i,j,k)$ are chosen as follows:
@@
Let $H_{2\text{-}D}$ denote the parity-check tensor constructed using the shifts defined in (8). Then, the corresponding Tanner graph is free of $4$-cycles, as stated in the following lemma.
@@
Consider arbitrary horizontal block-layer $k_{0}$ and $k_{1}$, where $0\leq k_{0},k_{1}\leq(p-1)$. To check the existence of cycle of length $4$, consider the following closed path of tensors indices:
@@
where $i_{s},j_{t}\in\{0,1,2,\ldots,(p-1)\}$ for all $s,t\in\{0,1\}$. Then, by Theorem 1, the $4$-cycle along the closed path $(i_{0},j_{0},k_{0}),(i_{1},j_{1},k_{0});(i_{1},j_{1},k_{1}),(i_{0},j_{0},k_{1})$ exists if and only if the following equations are satisfied under modulo $p$.
$a(i_{0},j_{0},k_{1})-a(i_{0},j_{0},k_{0})=a(i_{1},j_{1},k_{1})-a(i_{1},j_{1},k_{0}).$
$b(i_{0},j_{0},k_{1})-b(i_{0},j_{0},k_{0})=b(i_{1},j_{1},k_{1})-b(i_{1},j_{1},k_{0}).$
@@
where $\alpha=\left(\left\lfloor\frac{k_{1}}{p}\right\rfloor-\left\lfloor\frac{k_{0}}{p}\right\rfloor\right)\neq 0$. Therefore, for the existence of $4$-cycle we, must have
@@
The equations (9) and (10) cannot be satisfied simultaneously as $\phi$ and $\psi$ are bijective and $(i_{0},j_{0})\neq(i_{1},j_{1})$, proving the result. ∎
After establishing that the proposed parity-check tensor is free of $4$-cycles—and consequently that the associated two-dimensional code is also free of $4$-cycles—we next determine the exact rank of the proposed parity-check tensor, which in turn yields the exact code rate.

### v2

@@ Abstract
For any positive integer $g\geq 2$, we derive general condition for the existence of a $2g$-cycle in the Tanner graph of two-dimensional ($2$-D) classical quasi-cyclic (QC) low-density parity-check (LDPC) codes. Depending on whether $p$ is an odd prime or a composite number, we construct two distinct families of $2$-D classical QC-LDPC codes with girth >4>4 by stacking $p\times p\times p$ tensors. Furthermore, using generalized Behrend sequences, we propose an additional family of $2$-D classical QC-LDPC codes with girth >6>6, constructed via a similar tensor-stacking approach. All the proposed $2\text{-D}$ classical QC-LDPC codes exhibit an erasure correction capability of at least $p\times p$. Based on the constructed $2\text{-D}$ classical QC-LDPC codes, we derive two families of $2\text{-D}$ entanglement-assisted (EA) quantum low-density parity-check (QLDPC) codes. The first family of $2\text{-D}$ EA-QLDPC codes is obtained from a pair of $2\text{-D}$ classical QC-LDPC codes and is designed such that the unassisted part of the Tanner graph of the resulting EA-QLDPC code is free of $4$-cycles, while requiring only a single ebit to be shared across the quantum transceiver. The second family is constructed from a single $2\text{-D}$ classical QC-LDPC code whose Tanner graph is free from $4$-cycles. Moreover, the constructed EA-QLDPC codes inherit an erasure correction capability of $p\times p$, as the underlying classical codes possess the same erasure correction property.
@@ I Introduction
Following the success of one-dimensional (1-D) LDPC codes in classical communication and storage systems, higher-dimensional LDPC codes have also attracted significant research interest for the design of more powerful coding schemes. This growing interest is primarily driven by the stringent reliability requirements of advanced storage technologies such as shingled magnetic recording, two-dimensional (2-D) magnetic recording (TDMR) [16], and three-dimensional NAND flash memories [9], etc. In particular, a fundamental challenge in TDMR systems is that the dominant error mechanisms—including random errors, burst errors, and erasures—are inherently two-dimensional, arising from inter-track interference and inter-symbol interference. As a result, conventional $1$-D coding and decoding approaches are often inadequate and inefficient, motivating the development of native two-dimensional signal-processing and coding techniques attuned to the underlying channel structure [6, 3].
@@
2-D burst errors and erasures can be handled efficiently using $2$-D LDPC codes, which naturally extend the sparse graphical structure of classical LDPC codes to higher dimensions. In this direction, several works have investigated the construction, encoding, and decoding of $2$-D LDPC codes, demonstrating their effectiveness in mitigating structured $2\text{-D}$ error patterns [12, 8, 2]. These developments motivate the design of novel 2-D classical LDPC codes as a foundation for constructing quantum LDPC codes capable of correcting $2\text{-D}$ errors and erasures under Markovian constraints, which are particularly relevant for quantum storage and quantum communication systems [4, 7]. Although some studies have explored 2-D quantum LDPC code constructions, important limitations remain. For example, the authors of [17] construct quantum LDPC codes from 2-D spatially coupled LDPC codes which are not tailored to mitigate 2-D burst errors and erasures. Other recent works, such as [1] and [13], focus on quantum LDPC code design without explicitly addressing the correction of 2-D burst error/erasure patterns. Consequently, existing approaches do not fully account for correlated error structures arising from Markovian effects. This gap motivates our work, in which we first construct novel 2-D classical QC-LDPC codes and then utilize them to develop 2-D quantum LDPC codes specifically designed to correct 2-D errors and erasures under realistic Markovian noise models.
@@
In this work, we focus on the construction of $2\text{-D}$ entanglement-assisted (EA) quantum LDPC codes. We adopt the entanglement-assisted framework rather than the entanglement-unassisted setting to circumvent unavoidable cycles of length four that severely degrades the decoding performance. In other words, in the EA setting, the unassisted portion of the overall Tanner graph can be designed to be free of 4-cycles. Moreover, the entanglement-unassisted framework requires the underlying classical codes to satisfy the restrictive dual-containment constraint. In contrast, the availability of pre-shared ebits in the EA framework relaxes the stringent orthogonality requirements on the parity-check matrices, thereby allowing the construction of 2-D quantum LDPC codes from arbitrary classical codes. With the drive towards implementation of high-fidelity controlled quantum gates and caching of EPR pairs generated through a physical process, the EA construction of such 2-D QLDPC codes would be a promising direction. The main contributions of this work are summarized as follows.
@@
  1. (1) We first derive a general condition for the existence of $2g$-cycle in the Tanner graph of 2-D classical LDPC code. Depending on whether $p$ is an odd prime or a composite number, we construct two distinct families of $2$-D classical QC-LDPC codes with girth greater than $4$ by stacking $p\times p\times p$ tensors. Furthermore, using generalized Behrend sequences, we propose an additional family of $2$-D classical QC-LDPC codes with girth greater than $6$, constructed via a similar tensor-stacking approach. All the proposed code families achieve a burst erasure correction capability of at least $p\times p$.
@@
  2. (2) We then employ the constructed $2$-D classical QC-LDPC codes to develop two families of EA quantum LDPC codes. In the first family, two distinct $2$-D classical QC-LDPC codes are used to ensure that the unassisted portion of the overall Tanner graph is free of $4$-cycles, while requiring minimally a single ebit. The second family is constructed using a single $2$-D classical QC-LDPC code, where the underlying classical code itself is free of $4$-cycles.
@@
This article is organized as follows. In Section II, we first derive a general condition for the existence of $2g$-cycle in the Tanner graph of 2-D classical LDPC code. We then construct several families of 2-D classical QC-LDPC codes with burst erasure correction capabilities of at least $p\times p$. In Section III, we construct two families of 2-D EA quantum LDPC codes: one obtained from a pair of distinct $2$-D classical LDPC codes and the other constructed using a single $2$-D classical LDPC code. Finally, Section IV concludes the paper.
II 2-D Classical LDPC Codes Constructions
@@
From (3), it is evident that the application of $P$, $Q$ and $R$ to the identity tensor $I_{3\text{-D}}$ results in cyclic shifts along the positive $x$-, $y$-, and $z$-axes, respectively. Moreover, we refer to the shifts along the $x$ -, $y$-, and $z$-axes as the $X$-, $Y$-, and $Z$-shifts, respectively. These shifts have been illustrated for the identity tensor $I_{3\text{-D}}$ of size $3\times 3\times 3$ in Fig. 1.
@@
In light of Lemma 2, we construct 2-D classical LDPC codes by choosing permutations $P$ and $Q$ that implement circular shifts along the positive $x$- and $y$-axes, respectively. Based on this construction, we introduce the following assumptions and terminology to define the parity-check tensor $H_{2\text{-D}}$, which is obtained as a stack of tensors corresponding to the 2-D classical LDPC codes.
@@
    We introduce a new set of coordinate axes, denoted by $i$, $j$ and $k$, aligned with the $x$-, $y$-, and $z$-directions, respectively. The indices $(x,y,z)$ are used to label the positions of individual small cubes within the tensor, while the indices $(i,j,k)$ refer to the coordinates of a tensor as a whole. This distinction allows us to clearly differentiate between local indexing within the tensor and global indexing of the tensor in the parity-check tensor $H_{2\text{-D}}$.
@@
    For given size of identity tensor, i.e., $p\times p\times p$, we first define the X- and Y-shifts of a tensor located at an arbitrary position $(i,j,k)$ along the $x$- and $y$-directions. The X- and Y-shifts are denoted by $a(i,j,k)$ and $b(i,j,k)$, respectively.
@@
    For given positive integers $c$, $b$ and $h$ such that $p\mid h$, tensors with the prescribed shifts are stacked along the $i$-, $j$-, and $k$-directions. Within this stack, the $(i,j,k)$-th tensor, with indices satisfying $0\leq i\leq c$, $0\leq j\leq b$, and $0\leq k\leq h$, is constructed as
@@
To analyze the existence of cycles in the Tanner graph of $H_{2\text{-D}}$, we unfold the 3-D parity-check tensor $H_{2\text{-D}}$ into a $2\text{-D}$ parity-check matrix $H_{1\text{-D}}$ of size $ph\times bcp^{2}$. The unfolding is performed layer by layer along the vertical dimension: the first $cp$ columns of $H_{1\text{-D}}$ correspond to the first vertical layer of $H_{2\text{-D}}$, the next $cp$ columns correspond to the second vertical layer, and so on. Mathematically, the parity-check tensor $H_{2\text{-D}}$ is unfolded into the matrix $H_{1\text{-D}}$ such that the $(i,j)$-th entry of $H_{1\text{-D}}$ is the $(j\pmod{p},\left\lfloor\frac{j}{p}\right\rfloor,i)$-th entry in $H_{2\text{-D}}$.
Now, consider a closed path of length $2g$ in the parity-check tensor $H_{2\text{-D}}$, as described under assumption (9). For this closed path, let $\{\bar{k}_{i}\}_{i=0}^{g-1}$ denote the collection of $k$-indices, where each $\bar{k}_{i}$ represents the $k$-coordinate of a tensor located in the $k_{i}$-th horizontal block-layer and satisfies the conditions required to connect the corresponding cycle segments. Under these conditions, we arrive at the following theorem, which establishes the existence of cycles of length $2g$ in the Tanner graph associated with $H_{2\text{-D}}$ along the specified closed path.
@@
Let $H_{2\text{-D}}$ be the parity-check tensor composed of permutation tensors of size $p\times p\times p$. Assume that the permutation tensor located at position $(i,j,k)$ is specified by the shift functions $a(i,j,k)$ and $b(i,j,k)$. Consider the arbitrary closed path of length $2g$ as described under assumption (9), then the Tanner graph of $H_{2\text{-D}}$ contains a cycle of length $2g$ along the considered closed path if and only if there exists a collection of $k$-indices $\{\bar{k}_{i}\}_{i=0}^{g-1}$, where each $\bar{k}_{i}$ is the $k$-coordinate of a tensor located in the $k_{i}\text{-th}$ horizontal block-layer, satisfying the conditions required to connect the corresponding cycle segments, and
$\sum_{m=0}^{g-1}\Bigl(a(i_{m},j_{m},\bar{k}_{m})-a(i_{m+1},j_{m+1},\bar{k}_{m})\Bigr)$ $=0\pmod{p}$,
$\sum_{m=0}^{g-1}\Bigl(b(i_{m},j_{m},\bar{k}_{m})-b(i_{m+1},j_{m+1},\bar{k}_{m})\Bigr)$ $=0\pmod{p}$,
@@
Throughout the proof, $(i,j,k)$ denote the indices of the tensor, whereas $(x,y,z)$ denote the local indices of the small cubes, i.e., cubes within the tensor. Consider the arbitrary closed path of length $2g$ as follows:
@@
where $(i_{\ell},j_{\ell})\neq(i_{\ell+1},j_{\ell+1})$ and $k_{\ell}\neq k_{\ell+1}$ for $\ell=0,1,2,\ldots,g-2$, and $(i_{g-1},j_{g-1})\neq(i_{0},j_{0})$ with $k_{g-1}\neq k_{0}$. Next, consider the following sequence of local indices corresponding to the nonzero entries along the given closed path:
@@
where $0\leq x_{i},y_{i},z_{i}\leq p-1$ for all $i,j,k$, since without loss of generality these indices are taken to lie within the tensor, which is sufficient to prove our result. It is straightforward to observe that, for any tensor located at position $(i,j,k)$, the $x$ coordinate of a nonzero entry in the $z$-th layer of the tensor, where $0\leq z\leq p-1$, is given by
@@
Similarly, for any tensor located at position $(i,j,k)$, the $y$-coordinate of a nonzero entry in the $z$-th layer of the tensor, where $0\leq z\leq p-1$, is given by
@@
Thus, for the existence of a $2g$-cycle corresponding to the given closed path, it suffices to show that $(x_{g},y_{g})=(x_{0},y_{0})$. Here, $(x_{g},y_{g})$ denotes the $(x,y)$-coordinates of a nonzero entry in the layer $z=z_{g-1}$ of the tensor at position $(i_{0},j_{0},\bar{k}_{g-1})$, where $\bar{k}_{g-1}$ is the $k$-th coordinate of some tensor in the $k_{g-1}$ horizontal block-layer while $(x_{0},y_{0})$ denotes the $(x,y)$-coordinates of a nonzero entry in the layer $z=z_{0}$ of the tensor corresponding to $(i_{0},j_{0},\bar{k}_{0})$, where $\bar{k}_{0}$ is the $k$-th coordinate of some tensor in the $k_{0}$ horizontal block-layer. Now, by equation (4), for each index triplet along the closed path, the $x$-coordinates of the nonzero entries are given by the following set of equations:
$x_{0}$ $=z_{0}+a(i_{0},j_{0},\bar{k}_{0})\pmod{p}$,
$x_{1}$ $=z_{0}+a(i_{1},j_{1},\bar{k}_{0})\pmod{p}$,
$x_{1}$ $=z_{1}+a(i_{1},j_{1},\bar{k}_{1})\pmod{p}$,
$x_{2}$ $=z_{1}+a(i_{2},j_{2},\bar{k}_{1})\pmod{p}$,
@@
$x_{g-1}$ $=z_{g-1}+a(i_{g-1},j_{g-1},\bar{k}_{g-1})\pmod{p}$,
$x_{g}$ $=z_{g-1}+a(i_{0},j_{0},\bar{k}_{g-1})\pmod{p}$,
where each $\bar{k}_{i}$ is the $k$-coordinate of a tensor located in the $k_{i}\text{-th}$ horizontal block-layer, satisfying the conditions required to connect the corresponding cycle segments. By the above set of equations, we have the following equation under modulo $p$:
$x_{g}=x_{0}+\sum_{m=0}^{g-1}\Bigl(a(i_{m},j_{m},\bar{k}_{m})-a(i_{m+1},j_{m+1},\bar{k}_{m})\Bigr),$ (6)
where $(i_{g},j_{g})=(i_{0},j_{0})$. Now, by applying the same reasoning to the $y$-coordinates of the nonzero entries, we obtain the following equation under modulo p{p}:
$y_{g}=y_{0}+\sum_{m=0}^{g-1}\Bigl(b(i_{m},j_{m},\bar{k}_{m})-b(i_{m+1},j_{m+1},\bar{k}_{m})\Bigr),$ (7)
where $(i_{g},j_{g})=(i_{0},j_{0})$. For the existence of a $2g$-cycle, it is necessary that $(x_{g},y_{g})=(x_{0},y_{0})$. By equations (6) and (7), this occurs if and only if the following two equations are simultaneously satisfied:
$\sum_{m=0}^{g-1}\Bigl(a(i_{m},j_{m},\bar{k}_{m})-a(i_{m+1},j_{m+1},\bar{k}_{m})\Bigr)$ $=0\pmod{p}$,
$\sum_{m=0}^{g-1}\Bigl(b(i_{m},j_{m},\bar{k}_{m})-b(i_{m+1},j_{m+1},\bar{k}_{m})\Bigr)$ $=0\pmod{p}$,
@@
Henceforth, this section is divided into two subsections. The first subsection focuses on the construction of 2-D classical QC-LDPC codes with girth greater than $4$, while the second subsection addresses the construction of 2-D classical QC-LDPC codes with girth greater than $6$.
II-A 2-D Classical QC-LDPC Codes with Girth Greater than 44
Depending on whether $p$ is an odd prime or composite number, in this subsection, we construct two families of $2\text{-D}$ classical QC-LDPC codes by stacking permutation tensors of size $p\times p\times p$.
Let $p$ be an odd prime, and this assumption applies throughout this subsection unless stated otherwise. We construct the $2$-D QC-LDPC parity-check tensor $H_{2\text{-D}}$ by stacking $p\times p\times p^{2}$ permutation tensors, each of size $p\times p\times p$, along the $i$-, $j$-, and $k$-directions, respectively. These tensors are generated by applying various combinations of the shift operators $P$ and $Q$ to the identity tensor $I_{3\text{-D}}$, corresponding to $X$-shift and $Y$-shifts, respectively. While constructing $H_{2\text{-D}}$, the $(i,j,k)$th permutation tensor is given by $P^{a(i,j,k)}\circ Q^{b(i,j,k)}(I_{3\text{-D}})$, where $a(i,j,k)$ and $b(i,j,k)$ are chosen as follows:
@@
Let $H_{2\text{-D}}$ denote the parity-check tensor constructed using the shifts defined in (8). Then the corresponding Tanner graph is free of $4$-cycles as stated in the following lemma.
@@
Consider arbitrary horizontal block-layer $k_{0}$ and $k_{1}$, where $0\leq k_{0},k_{1}\leq(p-1)$. To check the existence of cycle of length $4$, consider the following arbitrary closed path of tensors indices:
@@
where $i_{s},j_{t}\in\{0,1,2,\ldots,(p-1)\}$ for all $s,t\in\{0,1\}$. Then, by Theorem 1, a $4$-cycle exists along the closed path $(i_{0},j_{0},k_{0}),(i_{1},j_{1},k_{0}),(i_{1},j_{1},k_{1}),(i_{0},j_{0},k_{1})$ if and only if there exists a set of $k$-indices $\{\bar{k}_{0},\bar{k}_{1}\}$, where each $\bar{k}_{i}$ is the $k$-coordinate of a tensor located in the $k_{i}\text{-th}$ horizontal block-layer, satisfying the conditions required to connect the corresponding cycle segments, and the following equations are simultaneously satisfied under modulo $p$:
$a(i_{0},j_{0},\bar{k}_{1})-a(i_{0},j_{0},\bar{k}_{0})=a(i_{1},j_{1},\bar{k}_{1})-a(i_{1},j_{1},\bar{k}_{0}),$
$b(i_{0},j_{0},\bar{k}_{1})-b(i_{0},j_{0},\bar{k}_{0})=b(i_{1},j_{1},\bar{k}_{1})-b(i_{1},j_{1},\bar{k}_{0}).$
@@
where $\alpha=\left(\left\lfloor\frac{\bar{k}_{1}}{p}\right\rfloor-\left\lfloor\frac{\bar{k}_{0}}{p}\right\rfloor\right)\neq 0$ $\pmod{p}$. Therefore, for the existence of $4$-cycle we, must have
@@
The equations (9) and (10) cannot be satisfied simultaneously as $\phi$ and $\psi$ are bijective and $(i_{0},j_{0})\neq(i_{1},j_{1})$. This completes the proof. ∎
After establishing that the proposed parity-check tensor is free of $4$-cycles—and consequently that the associated 2-D code is also free of $4$-cycles—we next determine the exact rank of the proposed parity-check tensor, which in turn yields the exact code rate.